OpenStudy (anonymous):

PLEASE HELP ASAP! The table below represents the distance of a car from its destination as a function of time. Part A: What is the y-intercept of the function and what does this tell you about the car? (4 points) Part B: Calculate the average rate of change of the function represented by the table between x = 1 to x = 3 and what does the average rate represent? (4 points) Part C: What would be the domain of this function if the car traveled the same rate until it reached its destination? (2 points)

OpenStudy (mathmale):

You'll need to post the table as well as the instructions. Please do that. Thanks.

OpenStudy (anonymous):

Time Distance (Hour) (miles) x y 0 900 1 850 2 800 3 750

OpenStudy (anonymous):

For part A I think the y-intercept is 50

undeadknight26 (undeadknight26):

I forgot how to do this stuff sorry.

OpenStudy (mathmale):

First: let's assume that the relationship between x and y is linear. Then take any two points from the table, for example: 1 850 2 800 and calculate the slope, using the slope formula. If youi've forgotten that, please see: http://www.purplemath.com/modules/slope.htm Once you've calculated the slope, type in your result here.

OpenStudy (anonymous):

I got 850 - 800 m = ————— 1 - 1 so 50?

OpenStudy (anonymous):

?

OpenStudy (mathmale):

Ask yourself: what would the denominator look like after you've simplified 1-1? Also, if you look carefully, you'll see that y decreases as x increases. Mind trying again to find the slope?

OpenStudy (anonymous):

-50?

OpenStudy (mathmale):

Great! Thanks for y our perseverance. Please be sure to use units of measurement: m = -50 miles / hour

OpenStudy (mathmale):

You are to find the y-intercept of the straight line through those two points. Use the slope-intercept formula. let your m (in that formula be -50 miles/hour. Pick either point and substitute its coordinates into the equation. If you're unsure of the point-slope formula, please say so (or better yet, look it up through an Internet search.

OpenStudy (anonymous):

Is this still for Part A?

OpenStudy (mathmale):

Yes, it definitely is about Part A. Your slope is correct. You could choose the point (1,850). The resulting point-slope form of the equation of the line is then \[y-y _{0}=m(x-x _{0})\] with \[x _{0}=1. and. y _{0}=850.\]

OpenStudy (mathmale):

Simplify the result; re-write the equation in the form y=m*x + b. " b " is the vertical intercept that you wanted.

OpenStudy (anonymous):

so y = -50x + b?

OpenStudy (anonymous):

the line crossed the 900 so is it y = -50x + 900?

OpenStudy (mathmale):

I haven't checked your result, but if it's correct, then yes, the y-intercept is b = 900, or, as a point, is (0,900). Unfortunately, I have to get off my computer now. if you still need help with these problems, perhaps I could help you later today.

OpenStudy (anonymous):

Are you able to answer Part B real quick?